Jaclyn can complete a job in 5.5 hours. Her brother Terry can complete the job in 4.5 hours. How long will it take them working together to complete the job?
Found 2 solutions by ptaylor, Alan3354:
Answer by ptaylor(2198)

Let x=amount of time it takes both working together to complete the job
Then together they work at the rate of 1/x of the job per hour
Jaclyn works at the rate of 1/(5.5) of the job per hour
Terry works at the rate of 1/(4.5) of the job per hour
So:
1/(5.5) + 1/(4.5)=1/x Multiply each term by (4.5)(5.5)(x)
4.5x+ 5.5x=24.75
10x=24.75
x=2.475 hours---amount of time it takes both working together to complete the job
CK
In 2.475 hours Jaclyn completes 2.475/5.5 of the job
In 2.475 hours Terry completes 2.475/4.5 of the job
2.475/5.5 + 2.475/4.5 should equal 1 (1 job that is)
0.45 + 0.55=1
1=1
Hope this helps---ptaylor

Answer by Alan3354(69443)

There's a shortcut to do these, product/sum
= 5.5*4.5/(5.5+4.5)
= 24.75/10
= 2.475 hours